# Agentic Visualization: Extracting Agent-based Design Patterns from Visualization Systems

### v1

@@ Abstract
Autonomous agents powered by Large Language Models are transforming AI, creating an imperative for the visualization field to embrace agentic frameworks. However, our field’s focus on a human in the sensemaking loop raises critical questions about autonomy, delegation, and coordination for such agentic visualization that preserve human agency while amplifying analytical capabilities. This paper addresses these questions by reinterpreting existing visualization systems with semi-automated or fully automatic AI components through an agentic lens. Based on this analysis, we extract a collection of design patterns for agentic visualization, including agentic roles, communication and coordination. These patterns provide a foundation for future agentic visualization systems that effectively harness AI agents while maintaining human insight and control.
@@
Autonomous agents—software entities that perceive their environment and take goal-directed actions—have been a cornerstone of artificial intelligence research for decades [22]. The emergence of Large Language Models (LLMs) has dramatically accelerated this paradigm, enabling general-purpose agentic AI systems capable of sophisticated planning, reasoning, and autonomous action in complex domains [20]. Given how previous waves of AI innovation have transformed visualization, this agentic turn will inevitably reshape our field. Yet visualization differs fundamentally from other AI application domains through its emphasis on human-centered reasoning, support for open-ended sensemaking, and placement of human analysts at the center of the analytical process. This creates an inherent tension between autonomous agent capabilities and visualization’s human-centered focus, raising questions about agency, delegation, coordination, and responsibility [6]. How can we design visualization systems that leverage agentic capabilities while preserving human insight and analytical control? What visualization tasks can we confidently delegate to agents?
To address these questions, we introduce the concept of agentic visualization (AV): interactive visual analysis systems that incorporate autonomous AI agents while maintaining human agency in the analytical process. Rather than starting from scratch, we analyze existing visualization systems with automatic or semi-automatic capabilities—sometimes described as mixed-initiative systems [7]—reinterpreting them as early forms of agentic visualization. By examining these systems through an agentic lens, we identify recurring structures that effectively coordinate automated components and human analysts. We formalize these structures as design patterns for agentic visualization that balance automation and analytical control. Design patterns, originating in architecture [1] and software engineering [5], provide a structured vocabulary for describing reusable solutions to common challenges. This approach aligns with human-centered AI principles that emphasize human values, interpretability, and agency in AI-augmented systems [2, 13].
@@
This paper makes the following contributions: (1) The concept of agentic visualization as a framework for designing and analyzing systems that incorporate autonomous agents while preserving human-centered analytical processes; (2) a collection of design patterns for agentic visualization extracted from a systematic analysis of prior literature on automated and semi-automated visualization systems; and (3) a demonstration showing how these patterns can be applied to a novel analytical scenario, including their composition into higher-order solutions. {mdframed}[backgroundcolor=blue!10, frametitle=
Sidebar: Agents and AI
, frametitlerule=true, frametitlebackgroundcolor=bgcolor]
Autonomous agents—software entities that perceive their environment and take goal-directed actions—have been a cornerstone of artificial intelligence research since the field’s inception in the 1950s [22]. The early reactive agents of the 1990s emphasized real-time interaction over complex reasoning, exemplified by Rodney Brooks’ subsumption architecture [17], followed by multi-agent systems and cognitive architectures that aimed to model distributed problem-solving and human cognition respectively [18].
@@
The current resurgence of agent technology, powered by Large Language Models (LLMs), has dramatically accelerated this paradigm, enabling general-purpose agentic AI systems capable of sophisticated planning, reasoning, and autonomous action in complex domains [20]. This represents a significant shift in capabilities but confronts persistent challenges that historical agent systems also faced: poor generalization across domains, scalability issues with complex tasks, inadequate coordination mechanisms, brittle performance in unexpected situations, and ethical concerns. These limitations highlight an inherent tension in agent design that was famously debated by Ben Shneiderman and Pattie Maes at IUI ’97 and CHI ’97, where Shneiderman advocated for direct manipulation interfaces that “affords the user control and predictability,” while Maes argued for delegation to software agents that could learn user preferences and act on their behalf [21]. This debate continues to shape the balance between computational agency and human control in interactive systems.
Addressing these challenges requires more than just technical improvements to agent capabilities. As Shah and White argue, successful agent systems need to generate sufficient value for users, provide adaptable personalization, establish trustworthiness, achieve social acceptability, and develop standardization protocols. Aridor and Lange’s work on design patterns for agent systems provides a useful foundation for capturing common solutions to recurring problems in agent design [16]. Their patterns are conceptually divided into three classes:
  • Traveling patterns: Agents moving and navigating in distributed environments;
  • Task patterns: Accomplishing agent objectives and distributing workloads; and
  • Interaction patterns: Governing communication and coordination between agents.
Recent work has extended this approach to foundation model-based agents, with researchers developing catalogs of architectural patterns that address challenges such as hallucinations, explainability, and accountability [19].
@@
The next step on that spectrum are agents that continuously work alongside the users without being explicitly triggered. They silently observe the users’ actions and offer unobtrusive help, possibly even extrapolating their needs and speculatively carrying out anticipated tasks ahead of time. This echoes ideas such as Mark Weiser’s Calm Computing, sometimes likened to “a raft of invisible servants” [11]. In visualization, we can find this idea in agent-based systems that run alongside or inside a visualization system and gently aid users, for example, in finding multi-view setups that maximize visibility and relevance of the shown data [9] or by suggesting meaningful next steps in visual analyses [14].
@@
AGENTIC VISUALIZATION
We define agentic visualization as the integration of autonomous AI agents with interactive visual analysis systems while balancing computational autonomy with human analytical control. This approach enables agents to independently perform complex analytical tasks while preserving the human analyst’s ability to guide the analytical process, interpret results, and maintain decision-making authority.
Agentic visualization extends and transforms several existing paradigms in visual analytics. Building on mixed-initiative interaction [7] and human-AI collaboration [2], agentic visualization specifically emphasizes agent autonomy beyond reactive assistance. Beyond merely responding to explicit user actions, agentic visualization allows for proactively identifying opportunities for analysis, suggesting approaches, or even pursuing analytical threads independently.
This shift challenges traditional interaction models where users issue commands and wait for responses. Instead, agentic visualization establishes a more fluid relationship where both human and agent can initiate analytical activities. This raises important questions about idle behavior: should an agent remain passive until directed (cold start), or should it proactively explore data and generate insights even without explicit user direction? The answer likely depends on context, task complexity, and user preferences.
But this evolution raises a critical question: how do we design agentic visualization systems that effectively balance agent autonomy with human control? This is particularly important in a field like visualization, which emphasizes the role of the human analyst in the interactive loop. Accordingly, the visualization field currently lacks systematic guidelines for determining appropriate levels of agent initiative, establishing effective communication between agents and analysts, and maintaining analytical transparency when agents operate with increasing independence.
To address these challenges, we look to past literature but reframe it through an agentic lens. By examining existing visualization systems with automatic or semi-automatic capabilities, we can identify patterns of effective human-agent collaboration that have emerged organically. These patterns, once formalized, can provide a foundation for designing the next generation of agentic visualization systems.
@@
To systematically identify patterns for agentic visualization, we conducted a comprehensive analysis of visualization literature where autonomous or semi-autonomous components collaborate with human analysts. Our methodology focused on extracting recurring design solutions that effectively balance computational agency with human control in visual analytics systems.
@@
    Communication Patterns: These patterns structure the information exchange between agents and human analysts. They address how agents present findings, receive feedback, and maintain transparency about their operations and reasoning processes.
@@
    Coordination Patterns: These patterns govern the temporal and workflow dynamics between agent and analyst activities. They establish when and how agents act, how concurrent human-agent actions are synchronized, and how analytical control shifts between participants.
@@
{mdframed}
[backgroundcolor=blue!10, frametitle=
Sidebar: Visualization and Agents
, frametitlerule=true, frametitlebackgroundcolor=bgcolor]
While agents have not been a major theme in the scientific community for visualization so far, there have been several instances of autonomous and semi-autonomous entities in visualization systems in the literature. Here we review these prior systems and their autonomous components.
  • InsightsFeed [23] is a progressive visual analytics system that generates and continuously refines an analysis of social media posts based on keywords, sentiment, and multidimensional embedding.
  • VisTrails [24] manages visualization workflows and their provenance metadata, enabling reproducibility, comparison of visualization pipelines, and systematic tracking of the scientific discovery process.
  • DataSite [25] is a proactive visual analytics system that autonomously executes background computations to identify salient data features and surfaces these findings as notifications in a feed timeline.
  • DataWeaver [26] is a data story authoring system that supports vis-to-text and text-to-vis compositions through statistics, custom algorithms, and LLMs.
  • SurfaceMapper [27] identifies relevant surfaces in 3D scientific data through multiple complementary agents that perform independent data interpretations.
  • HaLLMark [28] is a writing assistance system that autonomously captures and visualizes the complete provenance of writer-LLM interactions.
  • ProReveal [29] is a progressive visual analytics system that autonomously monitors intermediate analytical findings and activates safeguards when uncertain knowledge requires verification.
  • AVA [30] is a visualization agent framework that uses multimodal LLMs to autonomously interpret visual outputs and make visualization parameter decisions.
  • Medley [32] is a mixed-initiative system that recommends visualizations and widgets for dashboard composition, automatically providing interaction techniques.
  • StoryFacets [33] is a collaborative analysis and storytelling system that autonomously maintains visual provenance across multiple linked views.
  • Data Director [35] is a multi-agent system where LLM-powered agents generate insights, visualizations, and text to integrate them as a story into animated data videos.
  • Snowy [36] is a natural language interface that generates contextual utterance recommendations by combining data interestingness metrics and language pragmatics to guide analytic workflows.
  • DynaVis [37] is a multimodal system that combines a natural language interface with WIMP elements to generate standard charts and provide automatically synthesized widgets for visualization editing.
  • In-formation flocking [38] is a visualization approach that leverages a decentralized multi-agent system to recognize dynamic patterns and clusters in time-varying datasets, with local and global views.
  • InsightLens [39] is an LLM-agent-based interactive system that automatically captures and manages insights from users’ conversational data analysis workflow.
  • Voyager [40] generates galleries of visualization recommendations based on statistical and perceptual measures, enabling analysts to explore data without manual specification of variables or encodings.
@@
We identified four main agent roles for visualization.
@@
Pattern #1: The Forager pattern describe agents that look for data of interest according to parameters provided by the human. This pattern is associated with the challenges of processing and interpreting large datasets while keeping an overview of the dataset and adjusting to the domain context. Forager agents often work in groups, yet each Forager can either apply custom rules or variations of the given parameters to work efficiently in parallel and report back the results.
@@
*[regular] Examples: In SurfaceMapper [27], multiple agents with complementary expertise cooperate to identify data of interest in large volumes of scientific data. The agents report their findings to the human user and exchange them to revise their hypotheses. The in-formation flocking [38] technique involves multiple Forager agents that follow a set of behavior rules with the goal of highlighting potential patterns of interest in time-varying datasets. Each agent focuses on specific data neighborhoods and calculates similarity metrics to detect temporal changes. Working with Forager agents reduces the manual work of data processing and interpretation by collecting samples of interest and preliminary insights that the human can later compare and synthesize. Due to their flexibility, these agents can adjust to different domains. However, users still have to configure the agents according to their needs.
@@
Pattern #2: An Analyst agent generates and guides human users in making sense of a dataset. This pattern is associated with the challenges of exploring any data without specific goals to steer the process. To assist people who may struggle to choose a starting point for their exploration (i.e., the cold start problem), Analyst agents recommend multiple ways to start. The agent analyzes the dataset and prepares an initial set of recommendations. As soon as the person starts to interact, the agent starts learning from those actions to make follow-up recommendations. The Analyst encourages breadth-oriented exploration to increase data and analytical coverage.
@@
*[regular] Examples: Voyager [40] provides visualizations of all dataset variables to give an overview of the data space so the human user can choose what variables to focus on. In contrast, Snowy [36] generates multiple natural language inquiries based on patterns detected automatically in the dataset. The support of a computational Analyst allows users to have fewer concerns about leaving areas unexplored or missing key patterns. However, dealing constantly with updated recommendations may be overwhelming and distract the person from focusing on a specific data sample. Future work on Analyst agents should look for ways to balance these tradeoffs, adapting to the person’s skills.
@@
Pattern #3: Chart Creator agents assist or co-create visualizations with the human user. This pattern is associated with authoring scenarios, where the goal is to create and refine one or multiple visualizations. This involves a series of decisions, such as choosing an appropriate visual encoding and adjusting visualization parameters. When creating composite visualizations, such as a dashboard, choosing what visualizations to combine is also necessary. These choices are often made through multiple trials and repetitive actions. They require visualization literacy as well as technical knowledge, depending on the tool. Chart Creator agents help reduce the workload by helping the authoring process.
@@
*[regular] Examples: In DynaVis [37], the agent creates charts on demand and generates editing widgets when the person asks to change a visualization parameter. Medley [32] provides visualizations and widgets grouped by analytical intent that users can select from when authoring a visualization dashboard. The agent takes care of configuring standard interactions. One challenge of this pattern is that interacting with the agent is often the only way to create visualizations, potentially adding an extra layer of complexity for simple changes. Moreover, the automatically created elements are constrained to specific visualizations and widgets. Still, co-creating with an agent can reduce the need for repetitive tasks and advanced knowledge.
@@
Pattern #4: Storyteller patterns provide an agent who creates and updates data-driven stories. This pattern addresses the problem of generating and composing a cohesive story that brings together text, insights, and visual elements. Given that these processes require multiple skills regarding data analysis, text writing, and visualization authoring, the agent complements the skills of the human user by taking over one or multiple tasks. These agents help in the storytelling workflow by creating specific story elements.
@@
*[regular] Examples: DataWeaver [26] generates narrative text through dedicated algorithms and LLMs and integrates them into the main story. Data Director [35] provides an agent for designing and synchronizing visual animations for a data video. Designating storytelling tasks to agents helps people focus on specific story elements by delegating the generation or refinement of others. This can speed up the process and reduce barriers for people who may only have expertise in text writing, data wrangling, or visuals creation. The corresponding tradeoff is that the agent’s output needs to be checked for accuracy and appropriateness, which creates a new task for the human user and may represent a challenge if the person does not have the corresponding expertise to evaluate it.
@@
Pattern #5: The Insight Timeline pattern deals with situations where multiple agents, both humans and AI-powered ones, need a way to synchronize and share findings while continuing to work autonomously on their own tasks. The solution is to provide a common channel for agents and humans to share new insights and findings collected during the data analysis process. An Insight Timeline serves as a channel where both ongoing and explored tasks can be shared to (i) declare the current focus of exploration within the dataset to avoid redundancy, (ii) report interesting findings, and (iii) steer the analysis direction.
@@
*[regular] Examples: InsightsFeed [23] explores large textual datasets and progressively updates the results in an Insight Timeline so that the user can steer their analysis direction based on the results shared by the system. Similarly, DataSite [25] adopts the UI principles from InsightsFeed to propagate the results of background computation to the user without interrupting the cognitive process of the user. Users can also rank, sort, and filter the insights based on their needs.
@@
Pattern #6: The Progress Indicator pattern supports communication about which tasks the agents have undertaken and how far along each task is. This pattern helps human analysts interpret intermediate results, assess uncertainty in the current analysis, and identify areas of interest.
@@
*[regular] Examples: InsightsFeed [23] shows the progress and the aliveness by using a blinking rectangle next to a progress bar to highlight the current chunk of data being processed by the system. The system also uses a quality/stability metric to indicate the quality or uncertainty of the current state. They visualize the quality metric as a small line chart on top of each visualization to indicate if the quality of the computation is increasing, decreasing, or stable for the data behind the visualizations. ProReveal [29] shows the progression of both ongoing and completed visualizations to the analyst using a progress ring, which is hidden once the visualization is completely rendered with all the data. In both InsightsFeed and ProReveal, users have the option to interact with this process through pause, resume, and remove buttons to control the progression and explore the data in its current state. In addition to task progression, quality, and uncertainty, the progress indicators can also be used for other measures such as time spent vs. total time left, the amount of data processed, and the speed of the task performance.
@@
*[regular] Examples: VisTrails [24] captures provenance for visualization by tracking all the changes that users make to the visualization pipeline in an exploration task. Users can interact with the provenance information—presented as a directed graph—to reproduce a specific state and understand the process history. For collaborative exploration, VisTrails allows users to integrate other analysts’ workflows afterward in their view to analyze the provenance information. StoryFacets [33] presents a simpler view for displaying provenance from multiple collaborators on the same analysis session by creating multiple linked formats called facets. HaLLMark [28] uses summary statistics to encode the percentage of tasks performed by human and AI in a collaborative writing process.
@@
Agentic visualization workflows often employ multiple agents. Depending on the nature of the task, agents may work towards a shared goal, operate independently, or wait for others to complete their tasks before proceeding. Coordination is essential to ensure that agents and human analysts can schedule and execute their tasks effectively. Here we discuss different coordination patterns that describe how agents and human analysts coordinate during analysis.
@@
Pattern #8: At the beginning of the analysis process, it is often unclear which tasks should be prioritized to perform the analysis effectively, especially in large datasets. Scouting can serve as an initial coordination pattern, where multiple autonomous agents independently explore the dataset to identify potentially interesting patterns, anomalies, or other areas of interest. Human analysts can provide a single or multiple goals, and the agents can operate with this knowledge and adapt to the analysts’ feedback.
@@
*[regular] Examples: Mix&Match [10] releases smart particles (i.e., agents) into the dataset to highlight features of interest. The particles operate on rules, have a finite lifetime and scout for information in the dataset, leaving behind visual effects for the users. Users can also modify particle behavior by updating the rules to produce different effects. DataSite [25] proactively runs background computations on dataset to support human analysts during the different stages of the visual data exploration process. The analyst still maintains agency over the exploration and is actively supported by the system to generate insights, creating a balanced dialogue between the analyst and the system (or agents) and utilizing their respective strengths.
@@
Pattern #10: Monitoring enables agents launched during the analysis process to track temporally streaming data and notify the analyst of any changes, without overwhelming the human analyst.

### v2

@@ Abstract
Autonomous agents powered by Large Language Models are transforming AI, creating an imperative for the visualization area. However, our field’s focus on a human in the sensemaking loop raises critical questions about autonomy, delegation, and coordination for such agentic visualization that preserve human agency while amplifying analytical capabilities. This paper addresses these questions by reinterpreting existing visualization systems with semi-automated or fully automatic AI components through an agentic lens. Based on this analysis, we extract a collection of design patterns for agentic visualization, including agentic roles, communication and coordination. These patterns provide a foundation for future agentic visualization systems that effectively harness AI agents while maintaining human insight and control.
@@
Agents—software entities that perceive their environment and take goal-directed actions—have been a cornerstone of artificial intelligence research for decades [22]. The emergence of Large Language Models (LLMs) has dramatically accelerated this paradigm, enabling general-purpose agentic AI systems capable of sophisticated planning, reasoning, and autonomous action in complex domains [20]. Given how previous waves of AI innovation have transformed visualization, this agentic turn will inevitably reshape our field. Yet visualization differs fundamentally from other AI application domains through its emphasis on human-centered reasoning, support for open-ended sensemaking, and placement of human analysts at the center of the analytical process. The growing capabilities of agentic systems challenge the human-centered focus of visualization, raising questions about the level of agency, delegation, coordination, and responsibility between humans and AI [6]. What design patterns can guide us in developing agentic systems that leverage agentic capabilities while preserving human insight and analytical control?
To address these questions, we introduce the concept of agentic visualization (AV): interactive visual analysis systems that incorporate autonomous agents (whether AI-driven or rule-based) while maintaining human agency in the analytical process. Rather than starting from scratch, we analyze existing visualization systems both old and new with automatic or semi-automatic capabilities—sometimes described as mixed-initiative systems [7]—reinterpreting them as early forms of agentic visualization. While new visualization systems use LLM-based agents, we intentionally focus on prior foundational literature (before the introduction of LLMs), to understand what we can learn from the past to inform our present design patterns for agentic visualization.
By examining these systems through an agentic lens, we identify recurring structures that effectively coordinate automated components and human analysts. We formalize these structures as design patterns for agentic visualization that balance automation and analytical control. Design patterns, originating in architecture [1] and software engineering [5], provide a structured vocabulary for describing reusable solutions to common challenges. This approach aligns with human-centered AI principles that emphasize human values, interpretability, and agency in AI-augmented systems [2, 13].
@@
This paper makes the following contributions: (1) The concept of agentic visualization as a framework for designing and analyzing systems that incorporate autonomous agents while preserving human-centered analytical processes; (2) a collection of design patterns for agentic visualization extracted from a systematic analysis of prior foundational literature on automated and semi-automated visualization systems; and (3) a demonstration showing how these patterns can be applied to a novel analytical scenario, including their composition into higher-order solutions.
AGENTS AND AGENTIC VISUALIZATION
Agents are software entities that are capable of perceiving their environment and acting upon it to achieve goals [22]. An agent’s capacity to make their own decisions and take intentional actions towards that goals, is referred to as agency, that lies on a spectrum from simple, reactive behavior to fully autonomous behavior.
We define agentic visualization as the integration of agents with interactive visual analysis systems while balancing computational autonomy with human analytical control. This approach enables agents to independently perform complex analytical tasks while preserving the human analyst’s ability to guide the analytical process, interpret results, and maintain decision-making authority.
Agentic visualization extends and transforms several existing paradigms in visual analytics. Building on mixed-initiative interaction [7] and human-AI collaboration [2], agentic visualization specifically emphasizes agent autonomy beyond reactive assistance. Unlike traditional mixed-initiative systems [15], where humans and agents collaboratively initiate tasks, agentic visualization allows agents to proactively identify opportunities for analysis, suggest approaches, or even pursue analytical threads independently.
Instead of waiting for humans to issue commands and wait for response, agentic visualization challenges this traditional model by establishing a more fluid relationship where both human and agent can initiate analytical activities. This shift prompts questions about how agents should behave when idle: should they remain passive until directed (cold start), or should it proactively explore data and generate insights even without explicit user direction? The answer likely depends on context, task complexity, and user preferences.
To design such an agentic visualization system, we turn to both old and new visualization systems and reframe them through an agentic lens. Through this process, we can identify patterns of effective human-agent collaboration that have emerged organically.
@@
Autonomous agents—software entities that perceive their environment and take goal-directed actions—have been a cornerstone of artificial intelligence research since the field’s inception in the 1950s [22]. The early reactive agents of the 1990s emphasized real-time interaction over complex reasoning, exemplified by Rodney Brooks’ subsumption architecture [17], followed by multi-agent systems and cognitive architectures that aimed to model distributed problem-solving and human cognition respectively [18]. The current resurgence of agent technology, powered by Large Language Models (LLMs), has dramatically accelerated this paradigm, enabling general-purpose agentic AI systems capable of sophisticated planning, reasoning, and autonomous action in complex domains [20]. This represents a significant shift in capabilities but confronts persistent challenges that historical agent systems also faced: poor generalization across domains, scalability issues with complex tasks, inadequate coordination mechanisms, brittle performance in unexpected situations, and ethical concerns. These limitations highlight an inherent tension in agent design that was famously debated by Ben Shneiderman and Pattie Maes at IUI ’97 and CHI ’97, where Shneiderman advocated for direct manipulation interfaces that “affords the user control and predictability,” while Maes argued for delegation to software agents that could learn user preferences and act on their behalf [21]. This debate continues to shape the balance between computational agency and human control in interactive systems. Addressing these challenges requires more than just technical improvements to agent capabilities. As Shah and White argue, successful agent systems need to generate sufficient value for users, provide adaptable personalization, establish trustworthiness, achieve social acceptability, and develop standardization protocols. Aridor and Lange’s work on design patterns for agent systems provides a useful foundation for capturing common solutions to recurring problems in agent design [16]. Their patterns are conceptually divided into three classes: Traveling patterns: Agents moving and navigating in distributed environments; Task patterns: Accomplishing agent objectives and distributing workloads; and Interaction patterns: Governing communication and coordination between agents. Recent work has extended this approach to foundation model-based agents, with researchers developing catalogs of architectural patterns that address challenges such as hallucinations, explainability, and accountability [19].
@@
The next step on that spectrum are agents that operate continuously alongside the users without being explicitly triggered. They silently observe the users’ actions and offer unobtrusive help, even anticipating user goals by extrapolating their intent and performing likely tasks. This echoes ideas such as Mark Weiser’s Calm Computing, sometimes likened to “a raft of invisible servants” [11]. In visualization, we can find this idea in agent-based systems that run alongside or inside a visualization system and gently aid users, for example, in finding multi-view setups that maximize visibility and relevance of the shown data [9] or by suggesting meaningful next steps in visual analyses [14].
@@
To systematically identify patterns for agentic visualization, we conducted a comprehensive analysis of visualization literature where autonomous or semi-autonomous components collaborate with humans and amongst themselves. Our methodology focused on extracting recurring design solutions that effectively balance computational agency with human control in visual analytics systems.
@@
    Communication Patterns: These patterns structure the information exchange between agents and humans. They address how agents present findings, receive feedback, and maintain transparency about their operations and reasoning processes.
@@
    Coordination Patterns: These patterns govern the temporal and workflow dynamics between agent and user activities. They establish when and how agents act, how concurrent human-agent actions are synchronized, and how analytical control shifts between participants.
@@
While agents have not been a major theme in the scientific community for visualization so far, there have been several instances of autonomous and semi-autonomous entities in visualization systems in the literature. Here we review these prior systems and their autonomous components. InsightsFeed [23] is a progressive visual analytics system that generates and continuously refines an analysis of social media posts based on keywords, sentiment, and multidimensional embedding. VisTrails [24] manages visualization workflows and their provenance metadata, enabling reproducibility, comparison of visualization pipelines, and systematic tracking of the scientific discovery process. DataSite [25] is a proactive visual analytics system that autonomously executes background computations to identify salient data features and surfaces these findings as notifications in a feed timeline. DataWeaver [26] is a data story authoring system that supports vis-to-text and text-to-vis compositions through statistics, custom algorithms, and LLMs. SurfaceMapper [27] identifies relevant surfaces in 3D scientific data through multiple complementary agents that perform independent data interpretations. HaLLMark [28] is a writing assistance system that autonomously captures and visualizes the complete provenance of writer-LLM interactions. ProReveal [29] is a progressive visual analytics system that autonomously monitors intermediate analytical findings and activates safeguards when uncertain knowledge requires verification. AVA [30] is a visualization agent framework that uses multimodal LLMs to autonomously interpret visual outputs and make visualization parameter decisions. Data Formulator [32] is an interactive visualization authoring tool that uses AI-agent to create a visualization from raw data based on user’s intent provided as a data concept. StoryFacets [33] is a collaborative analysis and storytelling system that autonomously maintains visual provenance across multiple linked views. Data Director [35] is a multi-agent system where LLM-powered agents generate insights, visualizations, and text to integrate them as a story into animated data videos. Snowy [36] is a natural language interface that generates contextual utterance recommendations by combining data interestingness metrics and language pragmatics to guide analytic workflows. DynaVis [37] is a multimodal system that combines a natural language interface with WIMP elements to generate standard charts and provide automatically synthesized widgets for visualization editing. In-formation flocking [38] is a visualization approach that leverages a decentralized multi-agent system to recognize dynamic patterns and clusters in time-varying datasets, with local and global views. InsightLens [39] is an LLM-agent-based interactive system that automatically captures and manages insights from users’ conversational data analysis workflow. Jupybara [40] is an LLM-multi-agent-based Jupyter Notebook extension that supports analysts during exploratory data analysis and storytelling via a natural language interface.
@@
We identified four main agent roles for visualization. Each pattern is designed to support human-agent collaboration, where the human may have varying levels and types of expertise. We highlight relevant human roles for each pattern using the data science role described by Crisan et al. [4], as illustrated in Figure 1. These roles help inform how users engage with agents, but our primary focus remains on the agent’s roles and their behavior.
@@
Pattern #2: An Analyst agent generates and guides human users, typically a data analyst, in making sense of a dataset. This pattern is associated with the challenges of exploring any data without specific goals to steer the process. To assist people who may struggle to choose a starting point for their exploration (i.e., the cold start problem), Analyst agents recommend multiple ways to start. The agent analyzes the dataset and prepares an initial set of recommendations. As soon as the person starts to interact, the agent starts learning from those actions to make follow-up recommendations. The Analyst encourages breadth-oriented exploration to increase data and analytical coverage.
@@
*[regular] Examples: Snowy [36] generates multiple natural language inquiries based on patterns detected automatically in the dataset. Jupybara [40], an LLM-based multi-agent AI assistant, helps the data analyst as they explore their data in a Jupyter Notebook.
With the support of a Analyst, users worry less about missing key patterns or unexplored areas. However, frequent recommendations can become overwhelming and hinder focus. Jupybara [40] addresses this by displaying limited details for clarity. Future work could build on this to reduce distraction and provide clear information to the user.
@@
Pattern #3: Chart Creator agents assist users, such as technical analysts, domain experts, and evangelists, who regularly create and refine visualizations as part of their workflows. This pattern applies to scenarios where the goal is to design one or more visualizations, individually or as part of a composite view, such as a dashboard. This process involves a series of decisions, such as choosing an appropriate visual encoding and adjusting visualization parameters. These decisions often require multiple iterations and a combination of visualization literacy and technical skills, depending on the tool. Chart Creator agents help reduce the workload by helping the authoring process.
@@
*[regular] Examples: In DynaVis [37], the agent creates charts on demand and generates editing widgets when the person asks to change a visualization parameter. Data Formulator [32] uses AI-agent to transform raw data to visualizations based on authors defined data concepts. One challenge of this pattern is that interacting with the agent is often the only way to create visualizations, potentially adding an extra layer of complexity for simple changes. Moreover, the automatically created elements are constrained to specific visualizations and widgets. Still, co-creating with an agent can reduce the need for repetitive tasks and advanced knowledge.
@@
Pattern #4: Storyteller pattern provides agents that support the creation and refinement of data-driven stories by combining text, insights, and visual elements into a cohesive narrative. These agents complement the work of user such as evangelists, analysts, and domain experts, by taking over the time-intensive tasks such as text writing, visualization authoring, video creation, and adding annotations.
@@
*[regular] Examples: DataWeaver [26] generates narrative text through dedicated algorithms and LLMs and integrates them into the main story. Data Director [35] provides an agent for designing and synchronizing visual animations for a data video. Delegating storytelling tasks to agents allows users to focus on specific story elements while offloading the generation of others. This can speed up the process and lower barriers for users who may not have the necessary expertise. However, this delegation introduces a tradeoff, where the agent’s output must still be reviewed for accuracy, which can be challenging without sufficient knowledge to evaluate it.
@@
Pattern #5: The Insight Timeline pattern deals with situations where humans and agents need a way to synchronize and share findings while continuing to work autonomously on their own tasks. The solution is to provide a common channel for them to share new insights and findings collected during the data analysis process. An Insight Timeline serves as a channel where both ongoing and explored tasks can be shared to (i) declare the current focus of exploration within the dataset to avoid redundancy, (ii) report interesting findings, and (iii) steer the analysis direction.
@@
*[regular] Examples: InsightsFeed [23] explores large textual datasets and progressively updates the results in an Insight Timeline so that the user, such as analyst and domain experts, can steer their analysis direction based on the results shared by the system. Similarly, DataSite [25] adopts the UI principles from InsightsFeed to propagate the results of background computation to the user without interrupting the cognitive process of the user. Users can also rank, sort, and filter the insights based on their needs.
@@
Pattern #6: The Progress Indicator pattern supports communication about which tasks the agents have undertaken and how far along each task is. This pattern helps users, such as analysts, engineers, domain experts interpret intermediate results, assess uncertainty in the current analysis, and identify areas of interest.
@@
*[regular] Examples: InsightsFeed [23] shows the progress and aliveness by using a blinking rectangle next to a progress bar to highlight the current chunk of data being processed by the system. The system also uses a quality/stability metric to indicate the quality or uncertainty of the current state with a small line chart. ProReveal [29] shows the progression of both ongoing and completed visualizations to the analyst using a progress ring, which is hidden once the visualization is completely rendered with all the data. In both InsightsFeed and ProReveal, users can interact with this process through pause, resume, and remove buttons to control the progression and explore the data in its current state. The progress indicators can also be used for other measures such as time spent vs. total time left, the amount of data processed, and the speed of the task performance.
@@
*[regular] Examples: VisTrails [24] captures provenance for visualization by tracking all the changes that users, especially data shapers and analysts, make to the visualization pipeline in an exploration task. Users can interact with the provenance information—presented as a directed graph—to reproduce a specific state and understand the process history. For collaborative exploration, VisTrails allows users to integrate other analysts’ workflows afterward in their view to analyze the provenance information. StoryFacets [33] presents a simpler view for displaying provenance from multiple collaborators on the same analysis session by creating multiple linked formats called facets. HaLLMark [28] uses summary statistics to encode the percentage of tasks performed by human and AI in a collaborative writing process.
@@
Agentic visualization workflows often employ multiple agents. Depending on the nature of the task, agents may work towards a shared goal, operate independently, or wait for others to complete their tasks before proceeding. Coordination is essential to ensure that agents and humans can schedule and execute their tasks effectively. Here we discuss different coordination patterns that describe how agents and users coordinate during analysis.
@@
Pattern #8: At the beginning of the analysis process, it is often unclear which tasks should be prioritized to perform the analysis effectively, especially in large datasets. Scouting can serve as an initial coordination pattern, where multiple autonomous agents independently explore the dataset to identify potentially interesting patterns, anomalies, or other areas of interest. Users , typically data shapers, stewards, or domain experts can provide a single or multiple goals, and the agents can operate with this knowledge and adapt to their feedback.
@@
*[regular] Examples: Mix&Match [10] releases smart particles (i.e., agents) into the dataset to highlight features of interest. The particles operate on rules, have a finite lifetime and scout for information in the dataset, leaving behind visual effects for the users. Users can also modify particle behavior by updating the rules to produce different effects. DataSite [25] proactively runs background computations on dataset to support users during the different stages of the visual data exploration process. The user still maintains agency over the exploration and is actively supported by the system to generate insights, creating a balanced dialogue between the analyst and the system (or agents) and utilizing their respective strengths.
@@
Pattern #10: Monitoring enables agents launched during the analysis process to track temporally streaming data and notify the users, such as analysts, domain experts, among others of any changes, without overwhelming them.